70feet of fencing material to make a pen.the length is to be 5feet longer than the width whats the width
Answer by nerdybill(7384)

70feet of fencing material to make a pen.the length is to be 5feet longer than the width whats the width
.
Let w = width
then
w+5 = length
.
2(w + w+5) = 70
2(2w+5) = 70
2w+5 = 35
2w = 30
w = 15 feet (width)